OpenStudy (tori040403):

HELP WILL MEDAL AND FAN!!!!!!! The cooking compartment of a microwave oven measures 15 inches long, 14 inches wide, and 9 inches high. What is the volume of the cooking compartment of this microwave oven?

OpenStudy (krazykricket):

V= L*W*H V=15*14*9 V=1890

OpenStudy (tori040403):

thanks!

OpenStudy (krazykricket):

Your Welcome :)

OpenStudy (tori040403):

:)